Find the period, range, and amplitude of the cosine function y = 3/2 cos t/2
Answer by lwsshak3(11628)

Find the period, range, and amplitude of the cosine function y = 3/2 cos t/2
**
Equation for cos function: y=Acos(Bx-C), A=amplitude, period=2π/B, phase shift=C/B
For given trig function: y = 3/2 cos t/2
Amplitude:3/2
Range:[-3/2, 3/2]
B=1/2
Period=2π/B=4π